# GestureGPT: Toward Zero-shot Free-form Hand Gesture Understanding with Large Language Model Agents

### v4

@@ 4.2. Experiment 2: Online Video Streaming on PC
This scenario is set when a user is watching online video on a PC monitor. Grounding gestures to the correct function is much more challenging in this scenario compared to the previous one. The video streaming interface contains a considerably broader range of functions, with numbers up to 66 in some tasks from the previous 18 functions. Moreover, many functions have similar semantic meanings, such as the ”vlog channel” button and the ”anime channel” button, making them difficult to distinguish solely through gestures. Furthermore, the interactive buttons and elements on the screen are much smaller than the smart appliances in the previous experiment, which reduces the performance of gaze-based function differentiation. The size of function elements ranges from $0.27$ to $20.71$ cm2 ($\text{MEAN}=2.73$, $\text{SD}=3.67$). By navigating through an interface rich in functions and semantic complexities, we intend to explore the boundary of our system’s performance and understand whether it can differentiate user intentions with only minor differences, which is essential for real-world applications.
4.2.1   Experiment Setting and Procedure
  • Experimental Platform and Gesture Data Collection - Our experimental framework utilizes Python and Selenium to interact with a video streaming platform, specifically targeting the website ”[China] From the Spring and Autumn Period to the Prosperous Tang Dynasty (Season 1, 12 Episodes)” on Bilibili [^2] https://www.bilibili.com/video/BV1sh411j7A4/. The platform automates video control operations via Selenium. User gestures are captured using a 1080P resolution webcam.
@@ 4.2.1   Experiment Setting and Procedure
    • Gaze Data: The Tobii Eye Tracker 5 is used to collect gaze data from participants.
    • Interaction History and External Context: Interaction history was extracted from the task sequence, while the external contexts were predefined.
  • Task Descriptions - Eight tasks were designed to simulate common operations performed while watching videos.
  • Participants - We recruited 16 participants from four local schools, compensating them at a rate of \$12 per hour. Their ages ranged from $18$ to $35$ years ($\text{MEAN}=26.875,\text{SD}=4.689$), comprising $10$ males and $6$ females.
@@
Detailed information about the experiment simulation interface, the list of devices and their functions, and the task list is provided in Appendix B.2.
4.2.2   Results Analysis
@@
Figure 9. Illustration of results from Experiment 2: Online Video Scenario, encompassing all 8 tasks (across the x-axis) and 4 settings (indicated by hue). For each task, the 4 settings are arranged from left to right as follows: Baseline, Only Gaze, Only History and External, and All. Each specific bar is further divided into Top 1 / Top 3 / Top 5 accuracy (Top 1, Top 3, Top 5 in gradient brown) and Negative outcomes (in cyan), totaling 48 data points.
@@
In this scenario, we evaluate GestureGPT under the same 4 context settings. The results are presented in the right part of Table 2 as well as Figure 9.
@@
The main conclusion that GestureGPT can effectively incorporate various context information to reason and predict user intentions remains consistent. Specifically, the Top-1 accuracy improved from 19.53% to 37.50% when all types of context are available, and also significantly outperforms each single context setting: +11.20% compared to only the history and external setting, and +11.72% compared to the only gaze setting. Both contexts contribute to the final performance, and the performance greatly drops if either of them is excluded. Compared with Experiment 1, where gaze information brings relatively more utility, these results further demonstrate that it requires joint reasoning over all available contexts to achieve the best performance under such a more complicated scenario.
@@
An intriguing observation from our results was that the baseline performance was superior to that observed in the home scenario, thereby necessitating a task-specific exploration. Tasks such as pausing, playing videos, and selecting the next video (Tasks 4, 5, and 8), which require less context, exhibited strong performance across all tested context settings. This phenomenon suggests that the agent’s inherent knowledge of video streaming interfaces makes it more probable to infer actions that are more often seen in such a video streaming interface. To some extent, the agent’s success in identifying these operations can be attributed more to an educated guess rather than a calculated match of gesture and context, indicating a form of intuition derived from the model’s extensive training data. This intuition helps our system’s performance in certain tasks, but also acts as bias in other tasks. This phenomenon is discussed further in Section 5.4.
@@
Another special case is Task 7, which is semantic specific task, highlights the essential role of accurate gesture description in achieving precise gesture grounding. Out of 16 participants, 12 performed the ’thumbs-up’ gesture for the task. In such cases, if the agent accurately recognizes the gesture, the outcome is correct, notably reflecting in high Top-1 accuracy results. The failures predominantly occurred due to incorrect gesture segmentation. For example, because users formed a fist after lifting their hand, leading the agent to mistakenly focus on the action of making a fist, which leads to map the ”Full Screen” or ”volume” function. We conducted a pilot validation exercise on the gesture segmentation rule used in our system with a human labeler splitting gesture frames from the whole video on data from eight participants as the ground truth in gesture segmentation. Our results, compared to the ground truth, revealed a high recall rate of 95.28%, yet a precision of only 43.91%, reflecting the inclusion of non-gesture-related frames. Future enhancements could involve advanced object detection techniques to improve segmentation precision.
@@
In cases where gesture descriptions were poorly articulated, incorporating additional context proved beneficial. For example, considering interaction history context, such as users exiting full-screen mode, allowed the agent to infer that the user might want to perform operations unrelated to video control, like ”liking” a video. This various context adaptation significantly improved this task performance in Top-5 accuracy metrics in this task from 35.42% to 68.75%.
@@ 5. Discussion
5.1. Language Input vs. Visual Input for Agents
GestureGPT is currently built with a Large Language Model for its superior understanding and reasoning capability. For the Gesture Description Agent, an alternative implementation could employ a multimodal model, which seems more intuitive. So we tested replacing the rule-based Gesture Description Agent with GPT-4o’s vision in a video streaming scenario. With all context considered, the Top-1 and Top-5 accuracy results were 41% and 72%. These results demonstrate performance comparable to our rule-based module which is 37.5% and 73%, thereby validating the effectiveness of the rule-based approach.
@@
5.2. Gesture Description Quality Assessment
The quality of the gesture description agent was evaluated through an expert questionnaire. For this evaluation, we randomly sampled descriptions generated from three repetitions for 256 gestures across two scenarios, resulting in a compilation of 256 gestures and their descriptions. The same questionnaire was then distributed to two gesture experts, who rated each gesture description on a 5-point Likert scale. Both expert have published gesture-related research articles on premier conferences. One expert’s score was 3.28 (std=1.41), and the other’s was 3.74 (std=0.73). The positive ratings show that our description agent can capture key information of the gesture.
@@ 5.3. Context Selection in Complex Systems
Our experiments revealed that although adding more contexts is supposed to be informative and useful, for certain circumstances, it might also bring irrelevant noise and accordingly distract the reasoning process of LMM agents (See Top-5 accuracy of all setting vs only gaze setting in Home scenario). In the home scenario, most informative contexts are concentrated on gesture itself and gaze, while history and external provide relatively less contribution to the final performance.
On one hand, the inclusion of more heterogeneous contexts increases the complexity and difficulty of context organization and management, which then requires more competent agents to process and reason over them. On the other hand, contexts with less information entropy inevitably brings irrelevant implications, and may result in misleading of agents for specific cases. As a result, this underscores the importance of making informed trade-offs between context incorporation and agent capability to optimize the system performance.
@@ 6. Limitation and Future Work
6.1. Recognition of Non-Interactive Gestures
Currently, GestureGPT identifies meaningful gestures with the system triggered by the action of raising a hand above chest to start a gesture. However, this approach may lead to false activation, as many gestures not intended for interaction could inadvertently trigger the system. In real-world scenarios, where various non-interactive movements occur, this can interfere with the LLM model’s reasoning process, consequently reducing system performance.
There are already several studies addressing this issue. Researches have developed frameworks using body pose, gaze, and gesture data (Schwarz et al., 2014) or the multimodal feedback mechanism (Freeman et al., 2016) to predict interaction intent, while others have integrated EEG and gaze patterns to differentiate meaningful attention during tasks (Ge et al., 2023; Koochaki and Najafizadeh, 2018). Our system allows easy integration of advanced gaze pattern analysis algorithms and new input channels, which will greatly help in discerning the meaningful interactive gesture.
6.2. Integration and Utilization of Top-N Predictions
GestureGPT currently outputs a Top-5 candidate function list, which can be readily augmented with list selection interaction optimizations, as studied in many existing works. Specifically, Isomoto et al. (2022) developed a dwell selection system that employs machine learning to predict user intent solely based on eye movement data. Furthermore, statistical design of dynamic menus could also expedite user selection particularly in AR/VR contexts. G-Menu (Vanderdonckt and Petit, 2019) utilizes gestures and keywords to design a dynamic menu, aiding in swift function selection. These techniques and designs can be readily integrated into GestureGPT to achieve efficient user selection of the target function.
6.3. System Delay
@@
The overall time cost of the framework is predominantly occupied by the LLM-related costs, accounting for up to 90% of the current time expenditure. These costs include LLM API request costs. For each gesture understanding task, GestureGPT typically requires 1 request for the gesture description and 6 rounds of conversation between the Gesture Inference Agent and Context Management Agent, which indicates great impact of response delay on the realtimeness of our system. Without focusing on practicality, our conceptual framework averages 227 seconds and 38,785 tokens per task. Possible directions for improvements:
@@
  • KV-cache: API calls require concatenating long system prompts each round, exponentially increasing time costs. Our current implementation adds the entire context library (11,000 tokens) to each query. KV-cache saves computed attention values after each round, significantly reducing time. The task execution time drops to 20 seconds after removing the context from latter rounds.
  • Inference speed:Groq’s fast model inference services have an output speed of over 700 tokens per second. Combining Groq (without rate limits) and KV-cache, the first context round will take 2-3 seconds, while the remaining rounds taking a total of about 1-2 seconds (averaging 6 dialogue rounds with input of 200 tokens and output of 100 tokens each round). This would reduce the theoretical optimal latency to under 5 seconds.
@@
Even though we do not have access to KV-cache-ready high-performance LLMs nor computing acceleration hardware as used by Groq, we anticipate GestureGPT will attract interest from other research domains, fostering collaborative efforts towards a practical and accessible free-form gesture interface.

### v5

@@ 4.2. Experiment 2: Online Video Streaming on PC
This scenario is set when a user is watching online video on a PC monitor. Grounding gestures to the correct function is much more challenging in this scenario compared to the previous one. The video streaming interface contains a considerably broader range of functions, with numbers up to 66 in some tasks from the previous 18 functions. Moreover, many functions have similar semantic meanings, such as the “vlog channel” button and the “anime channel” button, making them difficult to distinguish solely through gestures. Furthermore, the interactive buttons and elements on the screen are much smaller than the smart appliances in the previous experiment, which reduces the performance of gaze-based function differentiation. The size of function elements ranges from $0.27$ to $20.71$ cm2 ($\text{MEAN}=2.73$, $\text{SD}=3.67$). By navigating through an interface rich in functions and semantic complexities, we intend to explore the boundary of our system’s performance and understand whether it can differentiate user intentions with only minor differences, which is essential for real-world applications.
4.2.1. Experiment Setting and Procedure
  • Experimental Platform and Gesture Data Collection - Our experimental framework utilizes Python and Selenium [^2] https://www.selenium.dev/ to interact with a video streaming platform, specifically targeting the website “[China] From the Spring and Autumn Period to the Prosperous Tang Dynasty (Season 1, 12 Episodes)” on Bilibili [^3] https://www.bilibili.com/video/BV1sh411j7A4/. The platform automates video control operations via Selenium. User gestures are captured using a 1080P resolution webcam.
@@ 4.2.1. Experiment Setting and Procedure
    • Gaze Information: The Tobii Eye Tracker 5 is used to collect gaze data from participants.
    • Interaction History and External Context Information: Interaction history was extracted from the task sequence, while the external contexts were predefined.
  • Task Descriptions - Eight tasks were designed to simulate common operations performed while watching videos (Table 8).
  • Participants - We recruited 16 participants from four local schools, compensating them at a rate of \$12 per hour. Their ages ranged from $18$ to $35$ years ($\text{MEAN}=26.875,\text{SD}=4.689$), comprising $10$ males and $6$ females. Participants included 5 undergraduate students, 8 graduate students, and 3 research engineers. None of the participants had prior experience with AR/VR devices, which helped minimize bias due to varying familiarity with such technology.
@@
4.2.2. Results Analysis
@@
Figure 9. Illustration of results from Experiment 2: Video Scenario, encompassing all 4 settings (across the x-axis) and 8 tasks (indicated by hue). For each setting, the 8 tasks are arranged from left to right as follows: task 1 Adjust Volumn, task 2 Video Progress Control, task 3 Enter Fullscreen, task 4 Pause Video, task 5 Play Video, task 6 Exit Fullscreen, task 7 Like the Video, and task 8 Next Video. Each specific bar is further divided into Top 1 / Top 3 / Top 5 accuracy (Top 1, Top 3, Top 5 in gradient brown) and Negative outcomes (in blue), totaling 48 data points.
@@
In this scenario, we evaluate GestureGPT under the same four context settings. The results are presented in the right part of Table 2 as well as Figure 9.
@@
The main conclusion that GestureGPT can effectively incorporate various context information to reason and predict user intentions remains consistent. Specifically, Top-1 accuracy improved from 19.53% to 37.50% when all types of context are available, significantly outperforming each single context setting: +11.20% compared to only the history and external setting, and +11.72% compared to the only gaze setting. Both contexts contribute to the final performance, and it greatly drops if either is excluded. Compared with Experiment 1, where gaze information brings relatively more utility, these results further demonstrate that it requires joint reasoning over all available contexts to achieve the best performance under such a more complicated scenario.
@@
An intriguing observation from our results was that the baseline performance was superior to that observed in the home scenario, thereby necessitating task-specific exploration. Tasks such as pausing, playing videos, and selecting the next video (Tasks 4, 5 and 8), which require less context, exhibited strong performance across all tested context settings. This phenomenon suggests that the agent’s inherent knowledge of video streaming interfaces makes it likely to infer actions commonly seen in such a video streaming interface. To some extent, the agent’s success in identifying these operations can be attributed more to an educated guess rather than a calculated match of gesture and context, indicating a form of intuition derived from the model’s extensive training data. This intuition helps our system’s performance in certain tasks, but also acts as bias in other tasks. This phenomenon is discussed further in Section 5.5.
@@
Another special case is Task 7, which is semantic specific task, highlights the essential role of accurate gesture description in achieving precise gesture grounding. Out of 16 participants, 12 performed the “thumbs-up” gesture for the task. In such cases, if the agent accurately recognizes the gesture, the outcome is correct, notably reflecting in high Top-1 accuracy results. The failures predominantly occurred due to incorrect gesture segmentation. For example, because users formed a fist after lifting their hand, leading the agent to mistakenly focus on the action of making a fist, which leads to map the “Full Screen” or “volume” function. We conducted a pilot validation exercise on the gesture segmentation rule used in our system with a human labeler splitting gesture frames from the whole video on data from eight participants as the ground truth in gesture segmentation. Our results, compared to the ground truth, revealed a high recall rate of 95.28%, yet a precision of only 43.91%, reflecting the inclusion of non-gesture-related frames. Future enhancements could involve advanced object detection techniques , differentiating interactive vs non-interactive gestures, or a specially designed beginning gesture to improve segmentation precision.
@@
In cases where gesture descriptions were poorly articulated, incorporating additional context proved beneficial. For example, considering interaction history context, such as users exiting full-screen mode, allowed the agent to infer that the user might want to perform operations unrelated to video control, like “liking” a video. This various context adaptation significantly improved this task performance in Top-5 accuracy metrics in this task from 35.42% to 68.75%.
@@
4.3. Result Analysis: Gesture Description Quality Assessment
The quality of the Gesture Description Agent was evaluated through an expert questionnaire. For this evaluation, we randomly sampled descriptions generated from three repetitions for 256 gestures across two scenarios, resulting in a compilation of 256 gestures and their descriptions. The same questionnaire was then distributed to two gesture experts, who rated each gesture description on a 5-point Likert scale (1 being “Almost no description is correct or relevant to the gesture” and 5 being “Almost all of the description is correct or relevant to the gesture”). Both expert have published gesture-related research articles on premier conferences. One expert scored 3.28 (std = 1.41), and the other scored 3.74 (std = 0.73). Their Kendall’s W score was 0.853, indicating a high level of agreement between the two experts. The positive ratings show that our description agent can capture key information of the gesture.
4.4. Pilot Study: Human Performance in Gesture Understanding
Figure 10. Pilot study interface showcasing. On the left, a video stream with a highlighted gaze trace indicating user focus during the task. On the right, a gesture recording from a video streaming scenario alongside a detailed interface function list. This setup was used to assess participant ability to map gestures to corresponding interface functions, facilitated by context such as gaze trajectories and historical information.
To gauge human performance on tasks similar to those managed by GestureGPT, we conducted a pilot study with four participants who were unfamiliar with the system and its experimental setups. We randomly selected data from four users involved in Smart Home and Video Streaming scenarios. During the pilot, participants were provided with video recordings that included the user’s gesture poses and movements, along with a list of interface functions available at that time. They were tasked with mapping each gesture to its corresponding function. The pilot study interface as shown in 10.
Initial attempts by participants were guided by gaze context, as demonstrated through gaze trajectory recordings, and historical context. Subsequent attempts were further supported by providing external context, the pilot study interface shown in 10. Subsequent attempts were further supported by providing external context. Results revealed significant challenges in gesture understanding even for humans: initial Top-1 and Top-5 accuracies were 43.8% and 81.3% for the smart home, and 18.8% and 50% for video streaming, respectively. With added external context, accuracies increased notably to 58.3% (Top-1) and 91.7% (Top-5) in the smart home, and to 83.3% (Top-1) and 100% (Top-5) in video streaming. This preliminary result underscores the inherent complexity of interpreting gestures, even for humans, in contextually rich settings. So compared with gesture recognition, gesture understanding is much more difficult and challenging.
@@ 5. Discussion
5.1. Integration and Utilization of Top-N Predictions
Due to the inherent complexity of gesture understanding—far surpassing that of mere gesture recognition—we approach this challenge as a recommendation task, which is why we show the Top-3 and Top-5 accuracies. GestureGPT currently outputs a Top-5 candidate function list, which can be readily augmented with list selection interaction optimizations, as studied in many existing works. Specifically, Isomoto et al. (2022) developed a dwell selection system that employs machine learning to predict user intent solely based on eye movement data. Furthermore, statistical design of dynamic menus could also expedite user selection particularly in AR/VR contexts. G-Menu (Vanderdonckt and Petit, 2019) utilizes gestures and keywords to design a dynamic menu, aiding in swift function selection. These techniques and designs can be readily integrated into GestureGPT to achieve efficient user selection of the target function. The rapid advancement of LLMs and their increasing capability as contextualized agents, LLM-driven frameworks like GestureGPT will provide improved performance so that its Top-1 accuracy reaches a directly applicable level, ultimately ease the burden of secondary selection of users.
5.2. Language Input (LLM) vs. Visual Input (LMM) for Agents
GestureGPT is currently built with a Large Language Model for its superior understanding and reasoning capability. For the Gesture Description Agent, an alternative implementation could employ a Large Multimodal Model (LMM), which seems more intuitive. So we tested replacing the rule-based Gesture Description Agent with GPT-4o’s vision in a video streaming scenario. With All Context considered, the Top-1 and Top-5 accuracy results were 41% and 72%. These results demonstrate performance comparable to our rule-based module which is 37.5% and 73%, thereby validating the effectiveness of the rule-based approach.
@@ 5.1. Integration and Utilization of Top-N Predictions
5.3. Single Agent vs. Multi Agent
Assuming privacy concerns regarding LMMs are not considered, employing gesture visuals with contextual information in a LMM-based single-agent architecture could significantly reduce the system delay. Unlike multi-agent systems, which require multiple reasoning cycles to interpret gestures, a single-agent approach can deliver immediate results within one iteration. We conducted a preliminary test with data from one user in the smart home scenario, and the LMM-based single-agent approach achieved a Top-1 accuracy of 25.0% and a Top-5 accuracy of 75%. Due to privacy concerns, comprehensive testing was limited because other users did not authorize the publication of data.
Although the end-to-end LMM approach exhibits shorter delays, it presents several disadvantages: it is cumbersome to adjust and train, lacks task transferability, and falls short in terms of scalability and generality. Conversely, the multi-agent architecture offers significant advantages, such as enabling component reuse, and facilitating modular training and optimization. Each architecture has its own applicable scenarios, and together, they can complement each other to achieve superior results.
However, the multi-agent architecture of GestureGPT inherently offers greater flexibility and is designed for broad applications. For example, instead of cameras, the description agent can work with sensors like IMUs, electromyography, and LiDAR, provided they can reconstruct the hand pose to some extent. The framework handles complex interaction contexts to mimic human gesture understanding, such as interpreting gestures linked to ongoing events. Integrating context into the inference agent would greatly increase the workload and limit flexibility. The independently implemented modules allow for easier optimization. Smaller models within the architecture are plug-and-play, thereby avoiding constant end-to-end training.
Furthermore, we contend that the multi-agent architecture of GestureGPT is ideally suited for complex interaction scenarios where single-step reasoning is insufficient. For example, an agent might need to interact with users or external databases before making decisions. Our architecture supports distributed processing, enabling different agents to handle specific tasks asynchronously, thereby having the potential to further enhance efficiency in managing complex interactions. For a detailed discussion on how this architecture addresses system delays, see Section 6.2.
@@ 5.4. Context Selection in Complex Systems
Our experiments revealed that although adding more contexts is supposed to be informative and useful, for certain circumstances, it might also bring overwhelming irrelevant information and accordingly distract the reasoning process of LMM agents (See Top-5 accuracy of All setting vs Only Gaze setting in Smart Home scenario). In Smart Home scenario, most informative contexts are concentrated on gesture itself and gaze, while history and external provide relatively less contribution to the final performance.
On one hand, the inclusion of more heterogeneous contexts increases the complexity and difficulty of context organization and management, which then requires more competent agents to process and reason over them. On the other hand, contexts with less information entropy inevitably brings irrelevant implications, and may result in misleading of agents for specific cases. This underscores the importance of making informed trade-offs between context incorporation and agent capability to optimize the system performance.
@@ 6. Limitation and Future Work
6.1. Handling Left-Handed Gestures, Multiple Hands, and Non-Interactive Gestures in Frame
GestureGPT currently identifies meaningful gestures when a right hand is raised above chest level. However, this implementation assumes that only one hand, typically the right hand, is visible. This assumption is often unrealistic, as interactions in real environments may involve left-handed users or require two-handed gestures, thereby limiting the system’s inclusivity. Furthermore, this simplification can lead to false activations, as unintended gestures or non-interactive movements may inadvertently trigger the system. This issue is exacerbated in scenarios where various non-interactive hand movements and multiple hands are present within the camera frame, potentially confusing the model and degrading system performance. Evaluating the system’s robustness when both hands are visible or when left-handed gestures are involved is essential for fully understanding its practical capabilities. Future iterations of our system will focus on distinguishing between intentional gestures and extraneous hand movements, even in the presence of multiple hands, and on improving left-hand gesture recognition to better accommodate left-handed users.
Several existing studies have addressed these challenges. Research (Just et al., 2006) has explored techniques for distinguishing between the static and dynamic aspects of hand movements, along with the challenges of recognizing one- and two-handed gestures. Similarly, vision-based interaction approaches for augmented reality (Szalavári and Gervautz, 1997) emphasize the importance of simplifying hand models and applying user intention recognition methods to achieve robust performance in complex scenarios. Additionally, frameworks that incorporate body pose, gaze direction, and multimodal feedback mechanisms have shown promise in predicting interaction intent (Schwarz et al., 2014; Freeman et al., 2016), while other studies have integrated EEG and gaze patterns to differentiate between meaningful and irrelevant gestures during tasks (Ge et al., 2023; Koochaki and Najafizadeh, 2018).
The architecture of GestureGPT is designed to be adaptable and scalable, enabling the integration of these advancements to further enhance its capabilities. Incorporating a left-hand recognition module within the Gesture Description Agent would allow the system to support both left-hand and two-handed gestures. Additionally, expanding the contextual input processed by the Context Management Agent could improve the system’s ability to distinguish non-interactive gestures, thus reducing the likelihood of false activations. In the future, we plan to explore these challenges further, including validating the accuracy of left-hand gesture recognition, providing comprehensive support for two-handed gestures, and refining the system’s activation mechanisms to enhance its practicality and robustness.
6.2. System Delay and Applicability
@@
The overall time cost of the framework is primarily driven by LLM-related processes, accounting for up to 90% of the total time expenditure. These costs include LLM API request costs. For each gesture understanding task, GestureGPT typically requires 1 request for the gesture description and 6 rounds of conversation between the Gesture Inference Agent and Context Management Agent, highlighting the significant impact of response delay on the real-time performance of our system. In its current conceptual form, without a focus on practicality, our framework averages 227 seconds and 38,785 tokens per task.
Although the conceptual framework of GestureGPT demonstrates promising potential for HCI applications, the current implementation faces significant challenges related to system delays, which affect its real-world applicability. Addressing these delays is essential for transitioning GestureGPT from a theoretical model to a practical tool capable of functioning in dynamic, real-time environments. Possible directions for improvements include:
@@
  • KV-cache: API calls require concatenating long system prompts in each round, exponentially increasing time costs. Our current implementation appends the entire context library (11,000 tokens) to each query. Implementing a KV-cache would reduce the need to reprocess identical contexts, potentially decreasing interaction time to approximately 20 seconds per task. This enhancement is expected to streamline operations by preserving computed attention values for recurring queries.
  • High-Performance Hardware Utilization: Integrating advanced computational resources, such as Groq’s hardware, could further reduce processing times. Groq’s fast model inference services offer an output speed of over 700 tokens per second.
@@
By combining Groq and KV-cache, the first context round would take 2-3 seconds, with subsequent rounds taking a total of about 1-2 seconds (averaging 6 dialogue rounds, each with 200 tokens input and 100 tokens output). This would reduce the theoretical optimal latency to under 5 seconds. Although we do not currently have access to KV-cache-enabled high-performance LLMs or the computing acceleration hardware provided by Groq, we anticipate that GestureGPT will attract interest from other research domains, fostering collaborative efforts towards developing a practical and accessible free-form gesture interface.
@@
While the current system delays may limit immediate real-world applications, the outlined advancements and potential collaborations offer a promising path toward minimizing these delays. Future iterations of our system, supported by ongoing research and technological improvements, could effectively reduce latency to levels suitable for dynamic, real-time user interactions. Such developments would not only address the practical limitations discussed but also pave the way for broader adoption and impact across various domains, ultimately fulfilling the potential of real-time, intuitive gesture-based interfaces.